A stained glass window in a museum is in the shape of a rectangle with a semicircle on top. The height of the rectangle part of the window is twice the base. If the base of the window is 12 feet, what is the area of the window ?


1
Expert's answer
2021-02-24T11:01:02-0500


Refer to the figure,

 The height of the rectangle part of the window is twice the base. 

So, "h=12\times 2\ ft=24\ ft"

Radius of semicircle part"=\frac{12}{2}=6\ ft"

Area of window "=" Area of rectangle "+" Area of Semi-Circle

Area of window "=" "hb+\frac{\pi r^2}{2}=12\times 24+\frac{1}{2}\times 3.14\times 6^2"

Area of window"\ = 344.52 \ ft^2"